OpenStudy (anonymous):

Can anyone go over this one as if they are explaining it to someone who has never seen fractions? Solve using the elimination method 2x/5+y/3=7 x/3-y/9=3 I'm not good with fractions but I do know some about the elimination method. PLEASE HELP! I'm so confused!

OpenStudy (loser66):

take out a paper, I walk you through the stuff, now, I solve the 1st fraction first \[\frac{2x}{5}+\frac{y}{3} =7\] do you know how to make it has the same denominator?

OpenStudy (anonymous):

multiply the first fraction by 3 and the second by 5. Those are the two smallest numbers to make it the smallest common denominator. Right?

OpenStudy (loser66):

good girl, tell me what you have, please

OpenStudy (anonymous):

\[\frac{ 6x }{ 15 }+\frac{ 5y }{ 15 }=7\]

OpenStudy (loser66):

yes, now combine them into ONLY one fraction, do you know how to do it?

OpenStudy (loser66):

don't worry, just do it. me here, I can steer you on the right way.

OpenStudy (anonymous):

\[\frac{ 6x+5y }{ 30 }\] or would it be over 15

OpenStudy (loser66):

over 15, re write, girl. and don't forget the = sign, you still have it, right? write out all of them, please

OpenStudy (anonymous):

\[\frac{ 6x+5y }{ ?15 }=7\]

OpenStudy (anonymous):

didn't mean to put that question mark in it

OpenStudy (loser66):

yes, now, multiply both sides by 15, what do you have?

OpenStudy (anonymous):

that is where fractions have always gotten confusing. I know how to do the =7 side, but how do I go about multiplying the fraction side? Just the top or the entire fraction?

OpenStudy (loser66):

\[\frac{6x+5y}{15}*15= 7*15\]

OpenStudy (loser66):

does it make sense to you?

OpenStudy (anonymous):

\[(15)\frac{ 6x+5y }{ 15 }=7(15)\] but would it end up equaling to \[\frac{ 90x+75y }{ 225 }=105\] ?

OpenStudy (anonymous):

or would the 225 stay as 15?

OpenStudy (loser66):

hold on, don't go on that way, it's wrong way, will get ticket !! stop!!

OpenStudy (loser66):

now, see mine.